OpenStudy (juan337492):

The base of a triangle exceeds the height by 8 centimeters. If the area is 90 square centimeters, find the length of the base and the height of the triangle. 1/2((x)(x+8)=90

OpenStudy (juan337492):

i know the answer is h=10 length of base=18 i just have trouble understanding why

OpenStudy (ajspeller):

x = base x+8 = height Solve the equation for x and then make the substitutions into the expressions above.

OpenStudy (ajspeller):

whoops .. i meant x+8 = base x = height

OpenStudy (juan337492):

what do i do with the 1/2 . what I am trying to say is how do I use 1/2 in 1/2((x)(x+8)=90

OpenStudy (ajspeller):

To find the area of a triangle the formula is one half the base times the height.

OpenStudy (ajspeller):

divide both sides by 1/2 to end up with x(x+8)=180 Then distribute the x to end up with x^2+8x-180=0 Then factor. If you need help factoring see the tutorial below. https://www.youtube.com/watch?v=HAi6Qv5Jiho Hope this helps